### v1

# Survival Modeling from Whole Slide Images via Patch-Level Graph Clustering and Mixture Density Experts

Survival Modeling from Whole Slide Images via Patch-Level Graph Clustering and Mixture Density Experts
Ardhendu Sekhar Indian Institute of Technology Bombay, Mumbai, India Vasu Soni Indian Institute of Technology Bombay, Mumbai, India Keshav Aske Indian Institute of Technology Bombay, Mumbai, India Garima Jain Indian Council of Medical Research, New Delhi, India Pranav Jeevan Indian Institute of Technology Bombay, Mumbai, India Amit Sethi Indian Institute of Technology Bombay, Mumbai, India
1 Abstract
We introduce a modular framework for predicting cancer-specific survival from whole slide pathology images (WSIs) that significantly improves upon the state-of-the-art accuracy. Our method integrating four key components. Firstly, to tackle large size of WSIs, we use dynamic patch selection via quantile-based thresholding for isolating prognostically informative tissue regions. Secondly, we use graph-guided k-means clustering to capture phenotype-level heterogeneity through spatial and morphological coherence. Thirdly, we use attention mechanisms that model both intra- and inter-cluster relationships to contextualize local features within global spatial relations between various types of tissue compartments. Finally, we use an expert-guided mixture density modeling for estimating complex survival distributions using Gaussian mixture models. The proposed model achieves a concordance index of $0.712\pm 0.028$ and Brier score of $0.254\pm 0.018$ on TCGA-KIRC (renal cancer), and a concordance index of $0.645\pm 0.017$ and Brier score of $0.281\pm 0.031$ on TCGA-LUAD (lung adenocarcinoma). These results are significantly better than the state-of-art and demonstrate predictive potential of the proposed method across diverse cancer types.
Keywords: Survival Analysis, Whole Slide Images (WSIs), Graph-Guided Clustering, Attention Mechanisms, Expert-Guided Modeling, Histopathology, Deep Learning, Cancer Prognosis
2 Introduction
Accurate survival prediction for cancer patients plays a vital role in personalized oncology, enabling clinicians to tailor treatment plans, adjust monitoring schedules, and allocate healthcare resources more effectively. In recent years, whole slide images (WSIs) — high-resolution digital scans of hematoxylin and eosin (H&E) stained pathology slides — have emerged as a valuable data modality for prognostic modeling. WSIs capture a wealth of histological information, including tumor architecture, stromal patterns, immune cell infiltration, and spatial interactions within the tumor microenvironment (TME), all of which are known to be correlated with patient outcomes.
Despite their potential, WSIs present significant challenges for computational analysis. A single slide may contain billions of pixels, making end-to-end processing computationally prohibitive. Moreover, acquiring detailed annotations at the cellular or regional level is expensive, time-consuming, and often infeasible at scale. As a result, most approaches rely on weakly supervised learning paradigms and patch-based representations, where each slide is divided into smaller tiles or patches that are processed independently or aggregated using pooling strategies.
While such methods have achieved considerable success in cancer classification and subtyping tasks, survival prediction introduces additional complexity. Unlike classification, which often hinges on localized discriminative features, survival analysis requires the joint modeling of long-range dependencies, subtle morphological cues, and interactions among spatially distributed components within the TME. Traditional statistical models such as the Cox proportional hazards model are limited in their capacity to capture non-linear and high-dimensional relationships inherent in WSIs. Deep learning approaches offer a more expressive alternative, but designing models that are both accurate and interpretable for survival estimation remains a persistent challenge.
To address these limitations, we propose a comprehensive modular framework that integrates four key components:
  1. 1.
    A dynamic quantile-based patch selection strategy that identifies prognostically informative tissue regions while reducing noise and computational burden;
  2. 2.
    A graph-guided k-means clustering technique to capture phenotype-level heterogeneity by grouping spatially coherent and morphologically similar patches;
  3. 3.
    An attention mechanism that incorporates both intra-cluster attention to model fine-grained interactions among patches within each phenotype cluster, and inter-cluster attention to capture high-level contextual relationships across different phenotype clusters;
  4. 4.
    An expert-guided mixture density modeling module, which models survival distributions using Gaussian mixture models.
Together, these components form a unified and comprehensive pipeline for interpretable, flexible, and clinically meaningful survival modeling. By bridging spatial reasoning, phenotype abstraction, and probabilistic outcome modeling, our framework offers a robust tool for enhancing prognosis in real-world cancer cohorts.
3 Related Work
Deploying weakly supervised learning has been pivotal in being able to use WSIs for deep survival analysis, primarily through multiple instance learning (MIL). Early methods, such as ABMIL[1], CLAM[2], and DSMIL[3], used attention-based pooling to identify informative patches for slide-level survival or classification. However, these treat patches independently, neglecting spatial relationships—critical for prognosis. To address this, graph-based methods, such as patch-GCN[4], leverage graph convolutional networks (GCNs) for modeling spatial dependencies. DeepAttnMISL[5] incorporates phenotype-level clustering and hierarchical graph transformer[6] to aggregate patch features across resolutions. TransMIL[7] uses self-attention but sacrifices spatial granularity due to memory constraints. HIPT [8] uses a hierarchical vision transformer for multi-resolution feature extraction, excelling in survival prediction tasks. PathoGen-X [9] aligns histopathology features with genomic data, using transformer-based translation to enhance survival prediction, even with limited paired data. Survival mixture density networks (SMDNs[10]) and SCMIL[11] model survival distributions as Gaussian mixtures, enabling survival curve estimation. However, these techniques lack mechanisms to disentangle latent subpopulations or leverage spatial phenotypes. Our expert-guided mixture density modeling builds on these ideas by combining phenotype-aware learning with expert-driven density estimation. Each expert works on a different set of input features and predicts survival outcomes for distinct latent subtypes, while a gating network dynamically assigns cohort-level weights, enhancing flexibility and interpretability through WSI-derived spatial phenotypes.
4 Methodology
We propose a framework for whole slide image (WSI) analysis to predict survival by combining four key components: a dynamic patch selection mechanism using quantile-based thresholding, graph-guided k-means clustering to group task-relevant patches, an attention mechanism to model local and global interactions, and an expert-guided mixture density modeling for survival prediction. These components enable the model to select relevant patches, form meaningful clusters, capture relationships within and across clusters, and learn individualized survival distributions. The entire pipeline is illustrated in Figure 1(a). Below, we describe each component with its mathematical formulations and implementation details.
4.1 Dynamic Patch Selection via Quantile-Based Thresholding
Our pipeline first divides each whole slide image (WSI) into non-overlapping patches of size 256×256 pixels. A tissue detection heuristic is applied to eliminate background or non-informative regions, retaining only tissue-containing patches for downstream analysis. Deep feature representations for the retained patches are then extracted using a histopathology foundational model based Vision Transformer (ViT) encoder[8] $\mathrm{F}(x)$, which has been pre-trained on a large-scale WSI dataset using self-supervised learning techniques[12]. This process yields a patch-level feature matrix $\mathrm{Feat}\in\mathbb{R}^{n\times d}$, where $n$ denotes the number of retained patches per WSI and $d$ represents the dimensionality of the ViT embeddings (e.g., 384).
Figure 1: (a) Architecture of the proposed Survival Modeling from Whole Slide Images via Patch-Level Graph Clustering and Mixture Density Experts framework. (b) Design of the proposed Expert-Guided Mixture Density Modeling architecture.
To identify task-relevant patches ($\mathrm{P}_{\mathrm{sel}}$) and task-irrelevant patches ($\mathrm{P}_{\mathrm{rem}}$), we employ a dynamic thresholding mechanism that adjusts based on the distribution of importance scores for each WSI. This approach improves upon static thresholding (e.g., a fixed cutoff of 0.25) by adapting to variability in score distributions across WSIs.
The patch selection module processes a patch feature matrix $X\in\mathbb{R}^{B\times\mathrm{N}\times d}$, where $B$ is the batch size (typically 1 for WSIs), $\mathrm{N}$ is the number of patches (e.g., up to 84,365 for a WSI), and $d$ is the feature dimension (e.g., 384 from a Vision Transformer). A sequence of linear transformations, GELU activations, and a sigmoid layer compute importance scores:
$\mathrm{logits}=\sigma\left(W_{2}\cdot\mathrm{GELU}(W_{1}\cdot X+b_{1})+b_{2}\right)$ (1)
where $\mathrm{logits}\in\mathbb{R}^{B\times\mathrm{N}\times 1}$, $W_{1}\in\mathbb{R}^{d\times h}$, $W_{2}\in\mathbb{R}^{h\times 1}$, $b_{1}$, $b_{2}$ are biases, $h=256$ is the hidden size, and $\sigma$ is the sigmoid function. The importance-weighted patch features are computed as $\mathrm{P}=X\odot\mathrm{logits}$, preserving the original dimensionality.
The adaptive threshold $\tau_{q}$ is set as the $q$-th quantile (default: $q=0.25$) of the importance scores for each WSI:
$\tau_{q}=\mathrm{quantile}(\mathrm{logits}_{b},q),\quad\mathrm{logits}_{b}\in\mathbb{R}^{\mathrm{N}}$ (2)
where $\mathrm{logits}_{b}$ is the squeezed logits for batch $b$. Patches are selected as:
$\mathrm{P}_{\mathrm{sel}}=\{X[:,i,:]\mid\mathrm{logits}_{b}[i]>\tau_{q}\}$ (3)
$\quad\mathrm{P}_{\mathrm{rem}}=\{X[:,i,:]\mid\mathrm{logits}_{b}[i]\leq\tau_{q}\}$ (4)
The indices $I_{\mathrm{sel}}$ and $I_{\mathrm{rem}}$ corresponding to $\mathrm{P}_{\mathrm{sel}}$ and $\mathrm{P}_{\mathrm{rem}}$ are used in subsequent processing, with $\mathrm{P}_{\mathrm{sel}}$ passed to the clustering module. The quantile $q$ is tunable via a validation set, selecting approximately the top $(1-q)\cdot 100\%$ of patches (e.g., 75% for $q=0.25$, or approximately 63,274 patches for a WSI with 84,365 patches) as task-relevant, adapting to each WSI’s score distribution.
4.2 Graph-Guided K-Means Clustering of Relevant Patches
The task-relevant patches ($\mathrm{P}_{\mathrm{sel}}\in\mathbb{R}^{m\times d}$, where $m$ is the number of selected patches) are grouped into clusters using k-means clustering on a $k$-nearest neighbors (k-NN) graph that integrates morphological and spatial similarities. This ensures clusters reflect both patch appearance (e.g., tumor vs. stroma patterns) and spatial proximity, capturing local structures relevant to survival prediction.
Given $\mathrm{P}_{\mathrm{sel}}$ and corresponding coordinates $\mathrm{coords}\in\mathbb{R}^{1\times m\times 2}$, we process each WSI as follows:
1. Normalization: Normalize features and coordinates to ensure comparable scales:
$X_{\mathrm{norm}}=\frac{\mathrm{P}_{\mathrm{sel}}-\mu_{X}}{\sigma_{X}+\epsilon},\quad\mathrm{coords}_{\mathrm{norm}}=\frac{\mathrm{coords}-\mu_{\mathrm{coords}}}{\sigma_{\mathrm{coords}}+\epsilon}$ (5)
where $\mu_{X}$, $\sigma_{X}$ are the mean and standard deviation of $\mathrm{P}_{\mathrm{sel}}$, $\mu_{\mathrm{coords}}$, $\sigma_{\mathrm{coords}}$ are similarly defined, and $\epsilon=10^{-6}$. Features are further L2-normalized:
$X_{\mathrm{norm}}=\frac{X_{\mathrm{norm}}}{\|X_{\mathrm{norm}}\|_{2}}$ (6)
2. Similarity Computation: Compute morphological and spatial similarities: - Morphological similarity via cosine similarity:
$S_{\mathrm{morph}}=X_{\mathrm{norm}}\cdot X_{\mathrm{norm}}^{T}$ (7)
where $S_{\mathrm{morph}}\in\mathbb{R}^{m\times m}$. - Spatial similarity via Euclidean distance with an exponential kernel:
$D=\mathrm{cdist}(\mathrm{coords}_{\mathrm{norm}},\mathrm{coords}_{\mathrm{norm}},p=2)$ (8)
$\quad S_{\mathrm{spatial}}=\exp\left(-\frac{D}{\sigma_{D}}\right)$ (9)
where $\sigma_{D}$ is the standard deviation of $D$ plus $\epsilon$. - Combine similarities using learnable weights $w_{\mathrm{morph}}$, $w_{\mathrm{spatial}}$ (initialized at 0.8 and 0.2, softmax-normalized):
$S=w_{\mathrm{morph}}\cdot S_{\mathrm{morph}}+w_{\mathrm{spatial}}\cdot S_{\mathrm{spatial}}$ (10)
3. k-NN Graph Construction: Select the top $k$ neighbors (e.g., $k=10$) per patch based on $S$, forming a sparse graph $G\in\mathbb{R}^{m\times k}$ with normalized weights:
$G_{i,j}=\frac{S_{i,j}}{\sum_{j\in\mathcal{N}_{k}(i)}S_{i,j}+\epsilon}$ (11)
4. K-Means Clustering: GPU-accelerated K-Means clustering (via cuML) is applied to the patch-level features $G$ to partition them into $C$ clusters, where $C$ is either specified or computed as $m/\mathrm{cluster\_size}$ (e.g., $\mathrm{cluster\_size}=64$). The clustering objective minimizes the within-cluster sum of squared distances:
$\underset{\{\mu_{i}\}_{i=1}^{C}}{\arg\min}\sum_{i=1}^{C}\sum_{x\in\mathrm{C}_{i}}\|x-\mu_{i}\|_{2}^{2}$ (12)
where $\mu_{i}$ denotes the centroid of cluster $\mathrm{C}_{i}$. After clustering, patches are sorted based on their assigned cluster labels, resulting in clusters $\mathrm{C}_{1},\mathrm{C}_{2},\ldots,\mathrm{C}_{C}$ and the corresponding coordinate groupings.
4.3 Attention Mechanisms
To model both local and global interactions in WSIs, we employ an attention mechanism that captures intra-cluster relationships among patches within each cluster and inter-cluster relationships across cluster representatives. This approach enhances the representation of the tumor microenvironment by modeling local cellular patterns and broader interactions, such as tumor-stroma or vascular relationships. The process can be described as follows:
1. Intra-Cluster Attention: - Input: Clusters $\{\mathrm{C}_{1},\mathrm{C}_{2},\ldots,\mathrm{C}_{C}\}$ of task-relevant patch features $\mathrm{P}_{\mathrm{sel}}\in\mathbb{R}^{m\times d}$, obtained from k-NN-based K-Means. - For each cluster $\mathrm{C}_{i}\in\mathbb{R}^{m_{i}\times d}$, apply Multi-Head Self-Attention (MHSA)[13] with $h=8$ heads to model local interactions among patches:
$\mathrm{C}_{i}^{\prime}=\mathrm{MHSA}(\mathrm{C}_{i})+\mathrm{C}_{i}$ (13)
where $\mathrm{MHSA}(\mathrm{C}_{i})=\mathrm{Concat}(\mathrm{head}_{1},\ldots,\mathrm{head}_{h})W^{O}$, and each head computes:
$\mathrm{head}_{j}=\mathrm{Attention}(Q_{j},K_{j},V_{j})=\mathrm{softmax}\left(\frac{Q_{j}K_{j}^{T}}{\sqrt{d/h}}\right)V_{j}$ (14)
with $Q_{j}=\mathrm{C}_{i}W_{j}^{Q}$, $K_{j}=\mathrm{C}_{i}W_{j}^{K}$, $V_{j}=\mathrm{C}_{i}W_{j}^{V}$, and $W_{j}^{Q},W_{j}^{K},W_{j}^{V}\in\mathbb{R}^{d\times(d/h)}$. A residual connection and layer normalization are applied:
$\mathrm{C}_{i}^{\prime}=\mathrm{LayerNorm}(\mathrm{C}_{i}+\mathrm{Dropout}(\mathrm{MHSA}(\mathrm{C}_{i})))$ (15)
2. Cluster Representative Extraction: - Compute a representative feature $R_{i}\in\mathbb{R}^{d}$ for each refined cluster $\mathrm{C}_{i}^{\prime}$ as the mean of its patch features:
$R_{i}=\frac{1}{m_{i}}\sum_{j\in\mathrm{C}_{i}^{\prime}}\mathrm{p}_{j}^{\prime}$ (16)
where $\mathrm{p}_{j}^{\prime}$ are the refined patch embeddings. - Output: A matrix of representatives $R\in\mathbb{R}^{C\times d}$.
3. Inter-Cluster Attention: - Apply MHSA[13] to $R$ to model relationships across clusters:
$R^{\prime}=\mathrm{MHSA}(R)+R$ (17)
followed by residual connection and layer normalization. This captures global interactions across different WSI regions.
4. Feature Integration: We first concatenate refined patch features from all clusters as follows:
$\widetilde{\mathrm{P}}=\mathrm{Concat}(\mathrm{C}_{1}^{\prime},\mathrm{C}_{2}^{\prime},\ldots,\mathrm{C}_{C}^{\prime})$ (18)
- Expand the global representation to match $\widetilde{\mathrm{P}}$’s sequence length:
$R^{\prime}_{\mathrm{expanded}}=$ $\mathrm{mean}(R^{\prime},\mathrm{dim}=1,\ \mathrm{keepdim=True})$
.$\mathrm{.expand}(-1,\ \widetilde{\mathrm{P}}.\mathrm{shape}[1],\ -1)$ (19)
We then combine local and global features:
$\widehat{\mathrm{P}}=\widetilde{\mathrm{P}}+R^{\prime}_{\mathrm{expanded}}$ (20)
We then concatenate task-irrelevant patches (if filtering is enabled) to form the final patch-level representation:
$\mathrm{P}_{\mathrm{final}}=\mathrm{Concat}(\widehat{\mathrm{P}},\mathrm{P}_{\mathrm{rem}})$ (21)
5. WSI-Level Aggregation: Compute the final WSI feature via attention-weighted aggregation as in AMIL[1]:
$\mathrm{z}_{\mathrm{WSI}}=\sum_{i=1}^{n}\alpha_{i}\widehat{\mathrm{P}}_{i}$ (22)
where $\alpha_{i}=\mathrm{softmax}(W_{a}\cdot\tanh(W_{h}\widehat{\mathrm{P}}_{i}^{T}))$ are attention weights, and $W_{a}$, $W_{h}$ are learnable parameters.
4.4 Expert-Guided Mixture Density Modeling
The mixture-of-experts (MoE) framework provides a principled approach to modeling complex distributions by decomposing the prediction task into a set of specialized submodels, or experts, each responsible for a distinct region of the input space. As detailed in Bishop’s Pattern Recognition and Machine Learning[14], the final output is a weighted combination of expert predictions, with weights governed by a learnable gating function. This design allows the model to dynamically adapt to heterogeneity in the data and allocate different computational roles to different experts.
Motivated by this, we propose an Expert-Guided Mixture Density Modeling framework for estimating individualized survival probability distributions from whole-slide image (WSI) features. The proposed module is depicted in Figure 1(b). Our approach builds upon recent advances in modeling survival outcomes using mixture density networks[10, 11]. The architecture includes a shared encoder, a gating module, and two expert networks. Each expert predicts a Gaussian mixture model (GMM) over a transformed time domain, enabling the model to flexibly capture multimodal survival behavior while preserving interpretability and parameter efficiency.
To flexibly and stably model non-negative survival times $t\in\mathbb{R}_{+}$, we apply a transformation to the time domain and represent the resulting variable using a Gaussian mixture model (GMM) for each expert:
$t=g(y)=\log(1+\exp(y)),\quad y=g^{-1}(t)$ (23)
$\quad\left|\frac{dy}{dt}\right|=\frac{e^{t}}{e^{t}-1}$ (24)
Each expert $e\in\{1,2\}$ models the transformed time variable $y$ using a $K$-component GMM based on the WSI feature vector $\mathrm{z}_{\mathrm{WSI}}$:
$\mathrm{PDF}(y\mid\mathrm{z}_{\mathrm{WSI}},e)=\sum_{i=1}^{K}\lambda_{i}^{(e)}(\mathrm{z}_{\mathrm{WSI}})\cdot\mathcal{N}(y\mid\mu_{i}^{(e)},\sigma_{i}^{(e)2}).$ (25)
The mixture weights $\lambda_{i}^{(e)}(\mathrm{z}_{\mathrm{WSI}})$, computed via an expert-specific neural network with softmax output, represent the GMM component probabilities for expert $e$. We incorporate cohort-level learnable vectors to parameterize the GMM components. Specifically, we introduce a shared mean vector $\mathbf{P}_{\mu}\in\mathbb{R}^{K}$ and a shared standard deviation vector $\mathbf{P}_{\sigma}\in\mathbb{R}^{K}$, which are transformed per expert via linear layers:
$\mu^{(e)}=\mathbf{W}_{\mu}^{(e)}\cdot\mathbf{P}_{\mu},\quad\sigma^{(e)}=\mathrm{softplus}(\mathbf{W}_{\sigma}^{(e)}\cdot\mathbf{P}_{\sigma}),$ (26)
allowing each expert to modulate global survival anchors to learn individualized risk distributions.
The proposed model estimates cohort-level survival outcomes through three key functions derived from the underlying Gaussian Mixture Model (GMM) over a transformed time domain: the Transformed Probability Density Function (TPDF), the Cumulative Death Probability (CDP), and the Survival Probability Function (SPF). The TPDF captures the likelihood of a death event occurring precisely at a given time and accounts for the Jacobian of the inverse transformation applied to ensure stability over the positive time domain. The CDP represents the cumulative probability that a death event has occurred by time $t$, effectively modeling the cumulative distribution function (CDF) over survival time. In contrast, the SPF quantifies the probability that a patient survives beyond time $t$, computed as one minus the CDP. Together, these functions allow the model to estimate both pointwise likelihood and cumulative survival behavior, providing a probabilistic foundation for learning from both censored and uncensored survival data. The Survival Functions are as follows:
1. Transformed Probability Density Function (TPDF):
$\mathrm{TPDF}(t\mid\mathrm{z}_{\mathrm{WSI}},e)=$ $\left|\frac{dy}{dt}\right|\cdot\sum_{i=1}^{K}\lambda_{i}^{(e)}(\mathrm{z}_{\mathrm{WSI}})\cdot$
$\mathcal{N}(g^{-1}(t)\mid\mu_{i}^{(e)},\sigma_{i}^{(e)2})$ (27)
2. Cumulative Death Probability (CDP):
$\mathrm{CDP}(t\mid\mathrm{z}_{\mathrm{WSI}},e)=\sum_{i=1}^{K}\lambda_{i}^{(e)}(\mathrm{z}_{\mathrm{WSI}})\cdot\Phi\left(\frac{g^{-1}(t)-\mu_{i}^{(e)}}{\sigma_{i}^{(e)}}\right)$ (28)
3. Survival Probability Function (SPF):
$\mathrm{SPF}(t\mid\mathrm{z}_{\mathrm{WSI}},e)=1-\mathrm{CDP}(t\mid\mathrm{z}_{\mathrm{WSI}},e).$ (29)
To combine the outputs from multiple experts, a gating network assigns soft probabilities over the two experts based on the WSI-level representation:
$G(\mathrm{z}_{\mathrm{WSI}})=\mathrm{softmax}(W_{g}\cdot\phi(\mathrm{z}_{\mathrm{WSI}}))\in\mathbb{R}^{2}$ (30)
where $\phi(\mathrm{z}_{\mathrm{WSI}})$ is the WSI-level representation from the shared encoder.
The final prediction is the weighted sum of expert outputs:
$\mathrm{TPDF}(t\mid\mathrm{z}_{\mathrm{WSI}})=\sum_{e=1}^{2}G_{e}(\mathrm{z}_{\mathrm{WSI}})\cdot\mathrm{TPDF}(t\mid\mathrm{z}_{\mathrm{WSI}},e)$ (31)
$\mathrm{SPF}(t\mid\mathrm{z}_{\mathrm{WSI}})=\sum_{e=1}^{2}G_{e}(\mathrm{z}_{\mathrm{WSI}})\cdot\mathrm{SPF}(t\mid\mathrm{z}_{\mathrm{WSI}},e)$ (32)
4.5 Training Objective
In survival analysis, censorship refers to cases where the event of interest (e.g., death or relapse) has not occurred for certain patients within the observed follow-up period. These censored observations are informative and must be handled carefully in the training objective to avoid biased learning.
Let $t_{d}$ denote the observed time and $c\in\{0,1\}$ be the censoring indicator, where $c=1$ indicates an uncensored (event occurred) sample and $c=0$ indicates a censored (event not yet occurred) sample. To jointly model both censored and uncensored samples, we adopt a negative log-likelihood (NLL) formulation based on the Transformed Probability Density Function (TPDF) and the Survival Probability Function (SPF):
$\mathrm{L}_{\mathrm{NLL}}=$ $-c\cdot\log(\mathrm{TPDF}(t_{d}\mid\mathrm{z}_{\mathrm{WSI}}))$
$-(1-c)\cdot\log(\mathrm{SPF}(t_{d}\mid\mathrm{z}_{\mathrm{WSI}})).$ (33)
For uncensored data ($c=1$), the model maximizes the likelihood of observing an event exactly at time $t_{d}$ using the TPDF. For censored data ($c=0$), the model maximizes the probability that the event has not occurred until time $t_{d}$, i.e., the survival probability.
This formulation ensures that censored samples are effectively used to shape the survival curve without making assumptions about the exact event time beyond the censoring point. This is critical in clinical datasets, where censoring is common and discarding such data would result in substantial information loss.
To further improve learning, we introduce two regularization terms:
1. Expert Diversity Loss:
$\mathrm{L}_{\mathrm{div}}=\|\mu^{(1)}-\mu^{(2)}\|_{2}^{2},$ (34)
2. Entropy Regularization on Gating:
$\mathrm{L}_{\mathrm{ent}}=-\sum_{e=1}^{2}G_{e}(\mathrm{z}_{\mathrm{WSI}})\log(G_{e}(\mathrm{z}_{\mathrm{WSI}})+\epsilon).$ (35)
The total loss combines these components:
$\mathrm{L}_{\mathrm{total}}=\mathrm{L}_{\mathrm{NLL}}+\lambda_{\mathrm{div}}\cdot\mathrm{L}_{\mathrm{div}}+\lambda_{\mathrm{ent}}\cdot\mathrm{L}_{\mathrm{ent}}.$ (36)
5 Experiment Details
5.1 WSI Dataset Details
To evaluate the capability of our proposed model, we conducted experiments on the publicly available The Cancer Genome Atlas (TCGA)[15] Whole Slide Image(WSI) datasets: Lung Adenocarcinoma (LUAD) and Kidney Renal Clear Cell Carcinoma (KIRC), comprising 459 and 509 whole-slide images (WSIs), respectively. Each slide was processed at 20× magnification and segmented into non-overlapping patches of size 256×256 pixels. Non-informative white regions were filtered out using a tissue detection heuristic. The mean number of patches per slide was approximately 12,150 for TCGA-LUAD and 14,300 for TCGA-KIRC.
5.2 Training Configuration
The training setup involved setting the quantile threshold for patch selection to 0.25. During k-NN graph construction, the top 10 nearest neighbors were considered for each patch. Patch-level graph features were subsequently partitioned into 64 clusters. Multi-head self-attention (MHSA) was applied within each cluster using 8 attention heads to capture local interactions. The Expert-Guided Mixture Density Modeling module incorporated two experts, each representing a Gaussian Mixture Model (GMM) with 100 components. The model was trained using the Adam optimizer with a learning rate of $2\times 10^{-4}$, weight decay of $1\times 10^{-3}$, and a dropout rate of 0.1. Training was performed over 20 epochs with a batch size of 1. To ensure robust performance estimation, 5-fold cross-validation was conducted across all datasets and model components.
5.3 Performance Metrics
To comprehensively assess model performance beyond traditional metrics, we adopted enhanced evaluation measures that capture both discrimination and calibration over time. While the standard concordance index (C-Index) [16] provides a global ranking of predicted risks, it is limited in temporal granularity. Therefore, we employed the time-dependent concordance (TDC), which assesses the proportion of correctly ranked patient pairs at multiple time points within a predefined interval $[0,\tau]$, offering a dynamic perspective on discriminative ability. In addition, we used the Brier Score (BS) to evaluate the calibration of predicted survival probabilities by measuring the mean squared error between predicted and observed outcomes. To account for prediction accuracy across the entire time horizon, we calculated the Integrated Brier Score (IBS), which integrates the BS over the interval $[0,\tau]$. Higher TDC and lower IBS values indicate superior model performance. All reported results are presented as mean ± standard deviation across validation folds.
6 Results
Our method advances WSI-based survival analysis by integrating dynamic quantile-based patch selection, graph-guided k-means clustering, and an Expert-Guided Mixture Density Modeling framework. Unlike methods that identify key patches using fixed scoring [1, 3, 2] or employ GCNs over adjacency-based graphs [4, 17], we construct a k-NN graph using both morphological and spatial similarities to capture contextually meaningful patch relationships. Compared to SCMIL [11], which applies local attention over fixed patch subsets, our method leverages graph-guided clustering followed by multi-head self-attention to dynamically model both intra- and inter-cluster interactions. This allows the network to flexibly attend to structurally and morphologically relevant regions, offering a more adaptive and context-aware attention mechanism that better captures critical survival-related patterns across the whole slide. Furthermore, our Expert-Guided Mixture Density Modeling module captures individualized, multimodal survival distributions with improved calibration and discrimination, resulting in superior performance on time-dependent concordance and integrated Brier score metrics compared to SCMIL and Transformer-based baselines [18, 7]. The experimental results are summarized in Table 1. Our proposed method achieves superior performance on both the TCGA-LUAD and TCGA-KIRC datasets, outperforming existing approaches in terms of both Time-Dependent Concordance (TDC) and Integrated Brier Score (IBS). Compared to prior methods, our approach consistently delivers the best results across both evaluation metrics.
Table 1: Performance comparison of different models on TCGA-KIRC and TCGA-LUAD datasets using Time-Dependent Concordance (TDC; higher is better) and Integrated Brier Score (IBS; lower is better). Best results are shown in bold.
Method TCGA-KIRC TCGA-LUAD TDC $\uparrow$ IBS $\downarrow$ TDC $\uparrow$ IBS $\downarrow$ AMIL [1] 0.628 ± 0.065 0.287 ± 0.013 0.614 ± 0.038 0.304 ± 0.037 CLAM [2] 0.666 ± 0.032 0.288 ± 0.029 0.595 ± 0.051 0.306 ± 0.026 DSMIL [3] 0.645 ± 0.031 0.288 ± 0.015 0.583 ± 0.065 0.321 ± 0.015 PatchGCN [4] 0.674 ± 0.049 0.279 ± 0.026 0.582 ± 0.055 0.307 ± 0.045 TransMIL [7] 0.632 ± 0.036 0.289 ± 0.016 0.515 ± 0.037 0.319 ± 0.029 HIPT [18] 0.635 ± 0.041 0.270 ± 0.021 0.540 ± 0.025 0.289 ± 0.068 HGT [6] 0.634 ± 0.058 0.269 ± 0.033 0.601 ± 0.042 0.289 ± 0.052 SCMIL [11] 0.688 ± 0.037 0.268 ± 0.021 0.622 ± 0.015 0.288 ± 0.060 Ours 0.712 ± 0.028 0.254 ± 0.018 0.645 ± 0.017 0.281 ± 0.031
6.1 Ablation Study
To comprehensively assess the contribution of individual components in our proposed survival modeling pipeline, we performed a set of ablation studies centered on patch selection and attention modeling. We first examined the role of dynamic patch filtering by varying the quantile threshold that determines which patches are retained for downstream processing. Our observations indicate that a threshold that is too permissive introduces noise from irrelevant or non-informative tissue regions, diluting the model’s ability to focus on prognostically meaningful patterns. This highlights the need for a carefully chosen filtering strategy to ensure that only the most informative patches are retained. In addition to patch selection, we investigated the impact of architectural elements designed to model complex tissue interactions. Specifically, we removed the dynamic patch filtering module while keeping other components intact, which resulted in a noticeable degradation in performance. This confirms that pre-selecting discriminative patches plays a critical role in enhancing the signal-to-noise ratio during representation learning. We also evaluated the effect of excluding the cluster-wise multi-head self-attention mechanism that operates on the graph-partitioned patch clusters. Eliminating this component significantly reduced the model’s effectiveness, suggesting that modeling both intra-cluster and inter-cluster relationships is essential for capturing fine-grained spatial and morphological dependencies within the slide. Overall, our ablation studies (see Table 2) reveal that each component—dynamic filtering and localized attention over clustered features—contributes significantly to the model’s capacity to extract survival-relevant information from whole slide images. Their synergistic integration is key to achieving robust and accurate survival predictions.
Table 2: Ablation Study Evaluating the Role of Quantile Threshold, Dynamic Filtering, and Cluster Attention
Method TCGA-KIRC TCGA-LUAD TDC $\uparrow$ IBS $\downarrow$ TDC $\uparrow$ IBS $\downarrow$ Quantile threshold = 0.5 0.701 ± 0.036 0.256 ± 0.018 0.634 ± 0.022 0.282 ± 0.013 Quantile threshold = 0.75 0.692 ± 0.056 0.256 ± 0.069 0.625 ± 0.041 0.282 ± 0.059 w/o dynamic filtering 0.683 ± 0.023 0.265 ± 0.012 0.616 ± 0.053 0.312 ± 0.032 w/o cluster attention 0.675 ± 0.016 0.267 ± 0.041 0.608 ± 0.012 0.313 ± 0.011
6.2 Interpretability
To evaluate the interpretability and clinical relevance of our proposed survival modeling framework, we conducted Kaplan–Meier (KM) survival analysis by stratifying patients into high-risk and low-risk groups based on the predicted survival scores generated by our model. This stratification was performed separately for the TCGA-KIRC and TCGA-LUAD cohorts, with the resulting KM plots illustrated in Figure 2. In both datasets, the survival trajectories of the high-risk(red) and low-risk(green) groups show a clear and meaningful separation over time, suggesting that our model captures biologically and prognostically relevant features. For the TCGA-LUAD cohort, the model achieved a statistically significant log-rank test p-value of 0.049, indicating that the difference in survival distributions between the two risk groups is non-random. In the TCGA-KIRC dataset, the stratification was even more discriminative, yielding a p-value of 0.030, thereby reinforcing the robustness of our model’s risk predictions. These results validate the model’s capacity to extract and encode morphologically meaningful patterns associated with patient outcomes. The distinct separation in KM curves underscores the effectiveness of our approach in generating clinically interpretable risk scores. By leveraging graph-guided clustering, dynamic patch selection, and cluster-level attention, the model is able to focus on spatially localized, yet prognostically significant, tissue regions. This design not only enhances predictive performance but also improves transparency, offering a practical pathway for integrating deep survival models into real-world clinical decision-making pipelines.
Figure 2: Kaplan–Meier survival curves for TCGA-LUAD and TCGA-KIRC, stratified by predicted risk. Statistically significant survival differences are observed between high- and low-risk groups.
7 Conclusion
In this study, we proposed a novel framework for survival analysis from whole slide images (WSIs), integrating dynamic quantile-based patch selection, graph-guided k-means clustering, and an Expert-Guided Mixture Density Modeling approach. Our approach effectively models both local and global tissue-level patterns by constructing patch-level graphs using spatial and morphological cues, and by applying multi-head self-attention over cluster-partitioned features. The incorporation of expert-guided mixture density Modeling allows for capturing complex survival distributions, leading to improved calibration and discriminative power. Comprehensive experiments on TCGA-LUAD and TCGA-KIRC datasets demonstrate the superiority of our method over existing approaches, as evidenced by consistent gains in time-dependent concordance and integrated Brier score. Ablation studies confirm the contribution of each module, while Kaplan–Meier survival curves illustrate the clinical relevance and interpretability of the predicted risk scores. For future work, we aim to extend our framework to multi-modal survival modeling by incorporating genomic, transcriptomic, or radiology data alongside WSIs. We also plan to explore domain generalization strategies to improve robustness across datasets from different institutions. Finally, integrating uncertainty quantification could enhance model reliability and facilitate its adoption in clinical decision-making pipelines.